# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v6

@@ 2 Series and generating functions in $\mathbb{C}$
The next approximation lemma is useful.
@@
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$,
$\big|s-\sum_{x\in Z}h(x)\big|\leq\varepsilon\,.$}
@@
$\big|s-\sum_{x\in Z}h(x)\big|\leq\big|s-\sum_{x\in Y}h(x)\big|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
@@
Two operations on series are grouping and product. A partition of a set $X$ is a set $P$ of nonempty and disjoint sets such that $\bigcup P=X$. If $X$ is at most countable, then so is $P$ and every set $Z\in P$. If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every $Z\in P$ the subseries $\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
@@
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and if $P$ is a partition of $X$, then $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
@@
$c=\sup(\{\sum_{x\in Z}|h(x)|\colon\;\text{$Z\subset X$ and $Z$ is finite}\})\ \ (<+\infty)\,.$}
Let $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3 and take for $i\in[n]$ finite sets
$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i})\ \ (\subset Z_{i})\,.$
We form the disjoint union $Z_{0}=Z_{1}^{\prime}\cup\dots\cup Z_{n}^{\prime}$ ($\subset X$). Then
${\textstyle\sum_{i=1}^{n}\big|\sum_{x\in Z_{i}}h(x)\big|}$ $\leq$ ${\textstyle\sum_{i=1}^{n}\big|\sum_{x\in Z_{i}}h(x)-\sum_{x\in Z_{i}^{\prime}}h(x)\big|\,+}$
$+$ ${\textstyle\sum_{x\in Z_{0}}|h(x)|\leq\sum_{i=1}^{n}2^{-i}+c\leq 1+c}$
and $S_{P}$ absolutely converges.
@@
We again use Lemma 2.3 and take for $i\in[n]$ finite sets
@@
Then $X_{0}$ is finite and $X^{\prime}\subset X_{0}$. Also, $Z_{i}^{\prime}\subset Z_{i}^{\prime\prime}$ for every $i\in[n]$. We have
$|s-t|$ $\leq$ ${\textstyle|s-\sum_{x\in X_{0}}h(x)|+\sum_{i=1}^{n}|\sum_{x\in Z_{i}^{\prime\prime}}h(x)-\sum_{x\in Z_{i}}h(x)|\,+}$
$+$ ${\textstyle|\sum_{i=1}^{n}\sum_{x\in Z_{i}}h(x)-t|\leq\frac{\varepsilon}{3}+\sum_{i=1}^{n}2^{-i}\cdot\frac{\varepsilon}{3}+\frac{\varepsilon}{3}\leq\varepsilon\,.}$
@@
Let $X_{1}$, $X_{2}$ be sets and $X_{1}\times X_{2}$ be their Cartesian product. If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, are series, then the series
$S_{1}\cdot S_{2}=\sum_{\langle x,\,y\rangle\in X_{1}\times X_{2}}h_{1}(x)h_{2}(y)$}
is the product of $S_{1}$ and $S_{2}$.
@@
If $S_{1}$ and $S_{2}$ absolutely converge and have sums $s_{1}$ and $s_{2}$, respectively, then $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s_{1}s_{2}$.
Proof. We show that $S_{1}\cdot S_{2}$ absolutely converges. Let $c\geq 0$ be such that for any finite sets $Y_{i}\subset X_{i}$, $i=1,2$, we have $\sum_{x\in Y_{i}}|h_{i}(x)|\leq c$. Let $A\subset X_{1}\times X_{2}$ be a finite set. We take any finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
@@
We show that the sum $s$ of $S_{1}\cdot S_{1}$ equals $s_{1}s_{2}$. It suffices to prove that for any $\varepsilon\in(0,1)$ we have $|s-s_{1}s_{2}|\leq\varepsilon$. So let $\varepsilon\in(0,1)$. Using Lemma 2.3 we take finite sets
@@
$Z=Y(S_{1}\cdot S_{2},\frac{\varepsilon}{4})\ \ (\subset X_{1}\times X_{2})\,.$}
We take finite sets $X_{1}^{\prime\prime}$ and $X_{2}^{\prime\prime}$ such that $X_{1}^{\prime}\subset X_{1}^{\prime\prime}\subset X_{1}$, $X_{2}^{\prime}\subset X_{2}^{\prime\prime}\subset X_{2}$ and $Z\subset X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}$. Then $|s-s_{1}s_{2}|$ is at most
${\textstyle\big|s-\sum_{\langle x,\,y\rangle\in X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}}h_{1}(x)h_{2}(y)\big|+\big|\sum_{x\in X_{1}^{\prime\prime}}h_{1}(x)\sum_{y\in X_{2}^{\prime\prime}}h_{2}(y)-s_{1}s_{2}\big|}$
${\textstyle\leq\frac{\varepsilon}{4}+|(s_{1}+t_{1})(s_{2}+t_{2})-s_{1}s_{2}|\,,}$
@@
Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i\in[k]$, be $k\in\mathbb{N}$ series. We define their product as the series
$\prod_{i=1}^{k}S_{i}=\sum_{\overline{x}\in\overline{X}}h_{1}(x_{1})h_{2}(x_{2})\dots h_{k}(x_{k})\,$,}
where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle\in X_{1}\times\dots\times X_{k}=\overline{X}$. Iterating Proposition 2.5 we obtain the following result.
@@
If series $S_{1}$, $\dots$, $S_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
@@
If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if their product $S_{1}\cdot S_{2}$ absolutely converges and has sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
@@
We prove absolute convergence of $S_{1}$. We take any $y\in X_{2}$ with $h_{2}(y)\neq 0$. Then for every finite set $X\subset X_{1}$ we have
$\sum_{x\in X}|h_{1}(x)|=\frac{1}{|h_{2}(y)|}\sum_{\langle x,y\rangle\in X\times\{y\}}|h_{1}(x)h_{2}(y)|\leq\frac{c}{|h_{2}(y)|}\,.$}
Thus $S_{1}$ absolutely converges. For $S_{2}$ we argue similarly. $\Box$
It is clear that the assumptions on $S_{1}$ and $S_{2}$ cannot be omitted.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$. We obtain a sufficient condition for lightness of the weight $h$ in terms of its values. We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk with nonzero weight and length at most $n$.
Definition 2.8 (slimness)
A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is slim if for every $n\in\mathbb{N}_{0}$ there is a constant $c=c(n)>0$ such that for every vertex $u\in V(h,n)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
@@
Proof. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be slim. We show by induction on $n\in\mathbb{N}_{0}$ that the series
$\sum_{w\in W(n)}h(w)$}
absolutely converges. For $n=0$ it is trivial because $W(0)=\{\langle 1\rangle\}$. Let $n>0$ and let $c>0$ be a constant such that
@@
for every finite set $X\subset W(n-1)$, every $u\in V(h,n-1)$ and every finite set $X^{\prime}\subset\mathbb{N}\setminus\{u\}$. Let $X\subset W(n)$ be finite. We may assume that all walks in $X$ have nonzero weights. We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex. Let $Y$ be the set of walks $w^{\prime}$. For $w^{\prime}\in Y$ let
@@
$\leq$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot c\leq c^{2}\,.}$
@@
as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. It follows that if $G(h)$ has finite degrees then $h$ is slim and hence light.
Definition 2.10 ($\xi$-convexity)
Let $\xi\in\mathbb{C}$. We call a light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ $\xi$-convex if for every vertex $u\in V(h)$ the sum
@@
We say that $h$ is convex if it is $1$-convex.
We show that the sums in the definition exist.
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be light and $u\in V(h)$. Then the series
@@
Proof. There is a walk $w_{0}=\langle u_{0},\dots,u_{n}\rangle$ in $G(h)$ such that $u_{0}=1$ and $u_{n}=u$. Let $c=h(w_{0})$ ($\neq 0$) and $W$ be the set of walks $\langle u_{0},\dots,u_{n},u_{n+1}\rangle$ in $K_{\mathbb{N}}$. Thus vertices $u_{0}$, $\dots$, $u_{n}$ are fixed and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$. The series $\sum_{w\in W}h(w)$ is a subseries of $\sum_{w\in W(n+1)}h(w)$ and therefore absolutely converges. By Proposition 2.2,
@@
We survey results on generating functions. We use univariate formal power series $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}$ with complex coefficients. They form an integral domain denoted by $\mathbb{C}[[x]]$. For $U$ and $V$ in $\mathbb{C}[[x]]$,
$U+V=\sum_{n\geq 0}(u_{n}+v_{n})x^{n}\,\text{ and }\,U\cdot V=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}u_{n}x^{n}$ with $u_{0}\neq 0$.
Let $\mathbb{C}_{1}=\{x\in\mathbb{C}\colon\;|x|=1\}$ and $\xi\in\mathbb{C}_{1}$. We define

### v8

@@ 2.2 Linear combination, grouping and product of series
This proposition easily extends to linear combinations with more terms. The next approximation lemma for series is useful.
@@
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$ we have
$|s-h(Z)|\leq\varepsilon\,.$
@@
$|s-h(Z)|\leq|s-h(Y)|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
@@
Two important operations on series are grouping and product. A partition of a set $X$ is a set $P$ of nonempty and disjoint sets such that $\bigcup P=X$. If $X$ is at most countable, then so is $P$ and every set $Z\in P$. If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every $Z\in P$ the subseries $S_{Z}=\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
@@
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and if $P$ is a partition of $X$, then the grouping series $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
@@
$c=\sup(\{\sum_{x\in Z}|h(x)|\colon\;\text{$Z\subset X$ and $Z$ is finite}\})\ \ (<+\infty)$}
and $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3, take for $i=1,2,\dots,n$ finite sets
$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i})\ \ (\subset Z_{i})$
and define the finite set $Z_{0}=Z_{1}^{\prime}\cup Z_{2}^{\prime}\cup\dots\cup Z_{n}^{\prime}$ ($\subset X$). Then
${\textstyle\sum_{i=1}^{n}|h(Z_{i})|}$ $\leq$ ${\textstyle\sum_{i=1}^{n}|h(Z_{i})-h(Z_{i}^{\prime})|+\sum_{x\in Z_{0}}|h(x)|}$
$\leq$ ${\textstyle\sum_{i=1}^{n}2^{-i}+c\leq 1+c}$
and we see that the series $S_{P}$ absolutely converges.
@@
We again use Lemma 2.3 and take for $i=1,2,\dots,n$ finite sets
@@
Then $X_{0}$ is a finite set and $X^{\prime}\subset X_{0}$. Also, $Z_{i}^{\prime}\subset Z_{i}^{\prime\prime}$ for every $i=1,2,\dots,n$. We have
$|s-t|$ $\leq$ ${\textstyle|s-h(X_{0})|+\sum_{i=1}^{n}|h(Z_{i}^{\prime\prime})-h(Z_{i})|+|\sum_{i=1}^{n}h(Z_{i})-t|}$
$\leq$ ${\textstyle\frac{\varepsilon}{3}+\sum_{i=1}^{n}2^{-i}\cdot\frac{\varepsilon}{3}+\frac{\varepsilon}{3}\leq\varepsilon\,.}$
@@
If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$ are series, then the series
$S_{1}\times S_{2}=\sum_{\langle x,\,y\rangle\in X_{1}\times X_{2}}h_{1}(x)\cdot h_{2}(y)$}
is called the product of the series $S_{1}$ and $S_{2}$.
@@
If $S_{1}$ and $S_{2}$ absolutely converge and have sums $s_{1}$ and $s_{2}$, respectively, then the product series $S_{1}\times S_{2}$ absolutely converges and has the sum $s_{1}s_{2}$.
Proof. We show that the series $S_{1}\times S_{2}$ absolutely converges. Let $c\geq 0$ be such that for any finite sets $Y_{i}\subset X_{i}$, $i=1,2$ we have $|h_{i}|(Y_{i})\leq c$. Let $A\subset X_{1}\times X_{2}$ be a finite set. We take finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
@@
We show that the sum $s$ of $S_{1}\times S_{2}$ equals $s_{1}s_{2}$. It suffices to prove that for every $\varepsilon\in(0,1)$ we have $|s-s_{1}s_{2}|\leq\varepsilon$. So let $\varepsilon\in(0,1)$. Using Lemma 2.3 we take finite sets
@@
$Z=Y(S_{1}\times S_{2},\frac{\varepsilon}{4})\ \ (\subset X_{1}\times X_{2})\,.$}
We take finite sets $X_{1}^{\prime\prime}$ and $X_{2}^{\prime\prime}$ such that $X_{1}^{\prime}\subset X_{1}^{\prime\prime}\subset X_{1}$, $X_{2}^{\prime}\subset X_{2}^{\prime\prime}\subset X_{2}$ and $Z\subset X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}$. Then
$|s-s_{1}s_{2}|$ $\leq$ ${\textstyle\big|s-\sum_{\langle x,\,y\rangle\in X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}}h_{1}(x)h_{2}(y)\big|+|h_{1}(X_{1}^{\prime\prime})h_{2}(X_{2}^{\prime\prime})-s_{1}s_{2}|}$
$\leq$ ${\textstyle\frac{\varepsilon}{4}+|(s_{1}+t_{1})(s_{2}+t_{2})-s_{1}s_{2}|\,,}$
@@
Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2,\dots,k$ be series. Their product is the series
$S_{1}\times S_{2}\times\dots\times S_{k}=\sum_{\overline{x}\in\overline{X}}h_{1}(x_{1})\cdot h_{2}(x_{2})\cdot\ldots\cdot h_{k}(x_{k})\,$,}
where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle$ and $\overline{X}=X_{1}\times\dots\times X_{k}$. By iterating Proposition 2.5 we get the following result.
@@
Let $S_{1}$, $S_{2}$, $\dots$, $S_{k}$ be absolutely convergent series with respective sums $s_{1}$, $s_{2}$, $\dots$, $s_{k}$. Then the series $S_{1}\times S_{2}\times\dots\times S_{k}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
@@
Let $S_{1}$ and $S_{2}$ be nonempty series that are not identically zero. If $S_{1}\times S_{2}$ absolutely converges and has sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge, and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
@@
We show that $S_{1}$ absolutely converges. Let $y_{0}\in X_{2}$ be such that $h_{2}(y_{0})\neq 0$. Then for every finite set $X\subset X_{1}$ we have
$|h_{1}|(X)=\frac{1}{|h_{2}(y_{0})|}\sum_{\langle x,y\rangle\in X\times\{y_{0}\}}|h_{1}(x)h_{2}(y)|\leq\frac{c}{|h_{2}(y_{0})|}\,.$}
Thus $S_{1}$ absolutely converges. For the series $S_{2}$ we argue similarly. $\Box$
It is clear that the assumptions on series $S_{1}$ and $S_{2}$ cannot be omitted.
2.3 Slim weights
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$. We obtain a sufficient condition for lightness of the weight $h$ in terms of its values. We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk with length at most $n$ and nonzero weight.
Definition 2.8 (slim weights)
A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is slim if for every $n$ in $\mathbb{N}_{0}$ there is a constant $c=c(n)>0$ such that for every vertex $u\in V(h,n)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
@@ 2.3 Slim weights
Proof. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be slim. We show by induction on $n\in\mathbb{N}_{0}$ that the series $\sum_{w\in D_{n}}h(w)$ absolutely converges. For $n=0$ it is trivial because $D_{0}=\{\langle 1\rangle\}$. Let $n>0$ and let $c>0$ be a constant such that
@@
for every finite set $X\subset D_{n-1}$, every $u\in V(h,n-1)$ and every finite set $X^{\prime}\subset\mathbb{N}\setminus\{u\}$.
Let $X\subset D_{n}$ be a finite set. We may assume that all walks in $X$ have nonzero weights. We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in D_{n-1}$ and $\ell(w)$ is the last vertex. Let $Y$ be the set of walks $w^{\prime}$. For $w^{\prime}\in Y$ let
@@
$\leq$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot c\leq c\cdot c=c^{2}\,.}$
@@
as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. If $G(h)$ has finite vertex degrees, then every series $\sum_{w\in D_{n}}h(w)$ has only finitely many nonzero summands and $h$ is therefore light.
2.4 $\xi$-convex weights
We generalize the concept of convexity.
Definition 2.10 ($\xi$-convex weights)
Suppose that $\xi\in\mathbb{C}$. We say that a light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $\xi$-convex if for every vertex $u\in V(h)$, the sum
@@ 2.4 $\xi$-convex weights
We say that a light weight is convex if it is $1$-convex.
Note that 0-convex weights are the ones that satisfy Kirchhoff’s second law. We show that the sums in the definition exist.
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and let $u\in V(h)$. Then the series
@@
Proof. There is a walk $w_{0}=\langle u_{0},\dots,u_{n}\rangle$ in $G(h)$ such that $u_{0}=1$ and $u_{n}=u$. Let $c=h(w_{0})$ ($\neq 0$) and $W$ be the set of walks $\langle u_{0},\dots,u_{n},u_{n+1}\rangle$ in $K_{\mathbb{N}}$. Thus, the vertices $u_{0}$, $\dots$, $u_{n}$ are fixed, and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$. The series $\sum_{w\in W}h(w)$ is a subseries of $\sum_{w\in D_{n+1}}h(w)$ and therefore absolutely converges. By Proposition 2.2, the series
@@
We proceed to generating functions.
2.5 Sums $U(1)_{1}$ and $U(1)_{2}$ for a generating function $U(x)$
We use univariate formal power series
$U=U(x)=\sum_{n\geq 0}u_{n}x^{n}$}
with complex coefficients. They form an integral domain denoted by $\mathbb{C}[[x]]$. For $U=\sum_{n\geq 0}u_{n}x^{n}$ and $V=\sum_{n\geq 0}v_{n}x^{n}$ in $\mathbb{C}[[x]]$,
$U+V=\sum_{n\geq 0}(u_{n}+v_{n})x^{n}\,\text{ and }\,U\cdot V=UV=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
Neutral elements are $0=0x^{0}+0x^{1}+0x^{2}+\dots$ and $1=1x^{0}+0x^{1}+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}u_{n}x^{n}$ with $u_{0}\neq 0$.
Let $\mathbb{C}_{1}=\{x\in\mathbb{C}\colon\;|x|=1\}$ and $\xi\in\mathbb{C}_{1}$. We define the set